How long will it take a 2350 W motor to lift a 330 kg piano to a six story window locate 16.0 m a above ground? Express your answer in seconds.
Answer #1

Solution)

Given,

Power, P=2350 W

Mass, m=330 kg

Height, h=16 m

We know,

Power, P= Force* height/time

Whete, Force, F=mg

So, time, T=mgh/P

Substitute values,

T=(330*9.8*16)/2350= 22.02 sec (Ans)

===========